# OM4OV: Leveraging Ontology Matching for Ontology Versioning

### v2

@@ VI Limitations
VII Conclusion
In this paper, we systematically analyse the similarities and differences between the OM and OV tasks and validate that they can share a unified pipeline with the necessary modifications. We propose a novel OM4OV pipeline with a cross-reference mechanism that leverages OM for OV. The new pipeline (1) overcomes several pitfalls in using OM for OV tasks, (2) significantly reduces the matching candidates, and (3) improves overall performance. We incorporate the OM4OV pipeline into a new OV system called Agent-OV, a functional extension of Agent-OM to handle OV tasks. Evaluations of three OAEI datasets validate the feasibility and reliability of our system. We also argue that the false mappings detected by OV systems are not necessarily actual false mappings.
Our approach is compatible with ontologies using different versioning methods (using URIs or additional versioning triples), and even ontologies missing or without version information. Our approach stores the version information independently from the ontology, offering a simple and lightweight way to track versioning changes in ontologies. In the future, we plan to apply the OM4OV pipeline in emerging domain ontologies that require consistent changes over time (e.g. smart building ontologies Brick Schema [23], RealEstateCore [24], and ASHRAE Standard 223P [25]).
Supplemental Material Statement
Agent-OV is an extended version of Agent-OM. Agent-OM can be downloaded from https://github.com/qzc438-research/ontology-llm. The OAEI track repositories used in this study can be downloaded from OAEI MELT at https://dwslab.github.io/melt/track-repository (date accessed: 2024-06-01). The OAEI data policy can be found in https://oaei.ontologymatching.org/doc/oaei-deontology.2.html.
Acknowledgment
The authors would like to thank Alice Richardson of the Statistical Support Network, Australian National University (ANU), for helpful advice on the statistical analysis in this paper. The authors also thank the Commonwealth Scientific and Industrial Research Organisation (CSIRO) for supporting this project.

### v3

@@ 7 Limitations
8 Conclusion
In this paper, we systematically analyse the similarities and differences between the OM and OV tasks and validate that they can share a unified pipeline with the necessary modifications. We propose a novel OM4OV pipeline with a cross-reference mechanism that leverages OM for OV. The new pipeline (1) overcomes several pitfalls in using OM for OV tasks, (2) significantly reduces the matching candidates, and (3) improves overall performance. We incorporate the OM4OV pipeline into a new OV system called Agent-OV, which serves as a functional extension of Agent-OM to handle OV tasks. Evaluations of our OAEI tested validate the feasibility and reliability of our system. Finally, we argue that the false mappings detected by OV systems are not necessarily actual false mappings.
Our approach is compatible with ontologies using different versioning methods (using URIs or additional versioning triples), and even ontologies missing or without version information. Our approach stores the version information independently from the ontology, offering a simple and lightweight way to track versioning changes in ontologies. In the future, we plan to apply the OM4OV pipeline in emerging domain ontologies that require consistent changes over time. For example, smart building ontologies Brick Schema [3], RealEstateCore [11], and ASHRAE Standard 223P [16].